OpenStudy (anonymous):

What is the compound inequality 8 < or = to 2x-4 < 2

OpenStudy (danjs):

hmm

OpenStudy (anonymous):

Hmm what?

OpenStudy (danjs):

\[8 \le (2x-4) \lt 2\] that right?

OpenStudy (anonymous):

Yes

OpenStudy (danjs):

that is a compound inequality

OpenStudy (danjs):

a<b<c

OpenStudy (anonymous):

Except there is no parentheses

OpenStudy (anonymous):

Ya but i have to find the solution to it

OpenStudy (danjs):

so you have 8 is less than or equal to 2x-4 AND 2x-4 is less than 2

OpenStudy (danjs):

oh, do you want to just isolate x to get the values for x, solve it

OpenStudy (anonymous):

My options are: 1. -6 < or = to X < -4 2. -2 < or = to X < -1 3. 0 < or = to X < 5 4. -2 < or = to X < 3

OpenStudy (danjs):

\[8 \le (2x-4) \lt 2\] \[8 +4\le (2x) \lt 2 + 4\] \[\frac{ 12 }{ 2 } \le x \lt \frac{ 6 }{ 2 }\] this says \[6 \le x \lt 3\] which is not true, x can not be bigger than 6 AND less than 3 at the same time

OpenStudy (danjs):

It ca be either OR

OpenStudy (danjs):

you mistyped something in the prob i think

OpenStudy (anonymous):

I'm confused

OpenStudy (danjs):

the way you put the problem, you cant have solutions for that, you can break it up into \[x \ge 6~~~~OR~~~~x \lt 3\]

OpenStudy (danjs):

a value can't be both those at the same time,

OpenStudy (anonymous):

I just put it the way that the school has it

OpenStudy (danjs):

you missing a negative sign, or something,

OpenStudy (danjs):

put the question down again using the draw tool

OpenStudy (danjs):

the inequality part

OpenStudy (anonymous):



OpenStudy (anonymous):

Is that what you mean @DanJS ?

OpenStudy (anonymous):

i say its 1

OpenStudy (anonymous):

can you explain why? I am so confused

OpenStudy (anonymous):

was it really \[-8 \le x-4\lt 2\]

OpenStudy (anonymous):

or \[-8 \le2 x-4\lt 2\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

add 4 all the way across\[-4\leq 2x<6\] then divide all by 2

OpenStudy (anonymous):

so it is \[-2lex <3\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[-2 \le x < 3\]

OpenStudy (anonymous):

\[-2\leq x<3\]

OpenStudy (anonymous):

Thank you so much!